# TerraCodec: Compressing Optical Earth Observation Data

### v2

@@ 1 Introduction
We address these gaps with TerraCodec, a family of neural codecs tailored to multispectral EO and pretrained on Sentinel-2 data. TerraCodec includes efficient image-based models for multispectral inputs: a lightweight Factorized Prior variant (TEC-FP) and an ELIC-based variant (TEC-ELIC) for optimal rate-distortion. We further introduce a Temporal Transformer (TEC-TT) that captures long-range dependencies across time without relying on hand-crafted motion priors. As shown in Fig. 1, these models offer a significantly better reconstruction than standard codecs at the same compression rate. At equal image quality, they reduce storage size by up to an order of magnitude, TEC-TT provides a further reduction for long time series by exploiting temporal structure.
@@ 3.1 TerraCodec
We introduce TerraCodec, a family of learned codecs for EO, including two image codecs: a lightweight factorized prior (TEC-FP) and a stronger space–channel context model (TEC-ELIC), as well as a temporal transformer (TEC-TT), with a flexible-rate variant (FlexTEC).
@@
Temporal Transformer (TEC-TT). TEC-TT builds on the VCT architecture [39]. We train a transformer to model temporal dependencies of seasonal EO data in latent space, predicting the current frame’s latent distribution from past context. Each frame $\mathbf{x}_{i}$ is encoded to latents $\mathbf{y}_{i}$ and quantized. We partition the current latent into $B$ non-overlapping spatial blocks $\{\hat{\mathbf{y}}_{i,b}\}_{b=1}^{B}$ and the two past latents into overlapping context blocks to increase the transformer’s receptive field. Each block $b$ is flattened into a sequence of $T$ tokens $\{\hat{y}_{i,b,t}\}_{t=1}^{T}$ of channel width $d_{\text{lat}}$. A temporal encoder aggregates the two previous frames into a joint context embedding $z_{\mathrm{joint}}=E(\hat{\mathbf{y}}_{i-2},\hat{\mathbf{y}}_{i-1})$. As shown in Figure 3, within each current block, a masked autoregressive transformer predicts token-wise prior parameters conditioned on already decoded tokens and $z_{\mathrm{joint}}$:
@@
We assume conditional independence across blocks given the context, allowing parallel probability estimation during encoding and parallel block decoding. Causal masking prevents attention to undecoded tokens; see the supplementary material for details. TEC-TT uses the same CNN analysis–synthesis transforms as TEC-ELIC. Unlike the original VCT, it is trained end-to-end on the rate–distortion objective without image pretraining, using a $\lambda$-schedule that emphasizes low-rate regimes early. We further adapt TEC-TT for flexible-rate scaling, introducing the FlexTEC variant in the next section.
@@ 3.2 Latent Repacking for flexible-rate models
Most neural codecs are trained for a fixed rate–distortion tradeoff. This makes deployment inflexible since achieving different bitrates requires retraining separate models. Our goal is to support variable rates at inference. We introduce Latent Repacking, which redistributes latent channels across tokens, and apply token masking with dynamic rate scaling during training so tokens learn an information-based ordering. Early tokens capture global structure, later ones refine detail. Truncating tokens then lowers bitrate while preserving global content. We demonstrate this by adapting TEC-TT, where strong priors allow missing tokens to be predicted, making it well-suited for Latent Repacking.
@@
In other words, $t^{\prime}_{u}$ is the $u$-th slice of $k$ consecutive channels of $A$, spanning the full spatial field. The procedure is reversible; reapplying slicing and repacking restores the original layout. After repacking, keeping the first $K$ tokens corresponds to $A[:,:,0:K\cdot k],$ the first $K\cdot k$ latent channels at every spatial location.
@@
We apply the approach on TEC-TT and introduce the Flexible-Rate TerraCodec (FlexTEC) model. The model applies Latent Repacking and masking inside the temporal transformer after image-wise compression and restores the original layout before image decoding (details in the supplementary material).
@@ 4.1 Pretraining
Temporal codec. TEC-TT is trained with a temporal context of two past frames for 300k steps with a global batch size of 24 on four NVIDIA A100 GPUs, requiring about 70 hours. We optimize with AdamW ($\text{lr}\!=\!10^{-4}$, weight decay $10^{-2}$) and employ half-cosine decay schedule to $\eta_{\min}=10^{-6}$ with 15% warmup steps. We sweep six $\lambda$ values for varying rate–distortion settings. For low-rate settings ($\lambda\leq 5.0$), we scale $\lambda$ by 10 during the first 15% of training before annealing back to the target value. Further pretraining details, including FlexTEC, are given in the supplementary material.
@@ 4.3 Downstream tasks
Cloud inpainting. TEC-TT captures spatiotemporal priors that can be applied beyond compression. We demonstrate this with zero-shot cloud removal on the AllClear benchmark [66]. Each sample consists of three cloudy observations with masks and one cloud-free target. To apply TEC-TT, we use the two least cloudy images as a temporal context and extract all cloud-free patches from the least cloudy image as input $x_{0}$. Patches in $x_{0}$ that are covered by clouds are predicted by TEC-TT. We report PSNR and other metrics, comparing against the official AllClear baselines. Following the benchmark, metrics are computed in auto mode across all bands, not per-band.
@@ 5.1 Rate–distortion
TEC-TT improves over the neural image models, although its margin over TEC-ELIC is smaller than the typical video–image gap. While EO sequences differ from natural video, being sampled at daily to seasonal rather than sub-second intervals, the short 4-frame validation setup further departs from typical video settings.
This limits temporal gains because half of the frames are bootstrap frames, which lack full previous context, whereas P-frames are predicted from two past frames. To better understand these effects, we study the role of temporal conditioning and report the RD performance on P-frames only in Figure 6. Without context, TEC-TT reduces to an image codec (labeled as image only). For a medium setting ($\lambda=5$), conditioning on one past frame improves compression by 13.6% compared to no context. On P-frames with two previous images, TEC-TT achieves a 22.6% rate reduction at equal PSNR, showing that longer EO sequences yield greater efficiency as bootstrap frames are amortized.
@@ 5.2 Flexible rate–distortion
Our TEC-TT-based FlexTEC model uses Latent Repacking to provide flexible rate-compression from a single checkpoint by transmitting a variable subset of latents and inferring the remainder from the model prior. Figure 8 compares FlexTEC against our fixed-rate models and standard codecs on P-frame compression. While fixed-rate TEC-TT models serve as an upper bound—each being optimally fitted for one specific rate–distortion—they require several separately trained models. In contrast, FlexTEC provides several user-controlled rate settings and performs close to or better than TEC-FP, depending on the setting. Further analysis in the supplementary material shows that FlexTEC encodes significant information in bootstrap frames, leading to similar bitrates independent of the token budget. The model then uses this information in the following P-frames to provide efficient rate–distortion settings. Figure 8 provides qualitative examples for different token budgets.
@@ 5.3 Downstream tasks
Model beliefs. Neural codecs rely on learned priors to estimate latent distributions for entropy coding. By decoding these priors into the image space, we obtain predictions that expose the model’s implicit knowledge. Figure 9 shows qualitative examples, with additional results in the supplementary material. We compare priors under three conditions: (a) no information from the current frame (only past context, TEC-TT), (b) partial information from the current latent (past frames and a subset of tokens for TEC-TT; subsets of channel groups for TEC-ELIC), and (c) full information of the latent. TEC-FP, lacking a hyperprior, cannot adapt to the latent at hand. Results show that TEC-TT already produces plausible forecasts from past frames alone and improves its beliefs when partial information becomes available, yielding the most refined predictions.
Figure 9: Model beliefs obtained by decoding learned priors into the image space. TEC-TT is shown with 0 (past context only), 5 and 16 tokens. TEC-ELIC uses limited and full channel-group context.
Cloud inpainting. Building on TEC-TT’s latent predictions, we evaluate zero-shot cloud removal on the AllClear benchmark [66]. TEC-TT is applied without task-specific training (see Sec. 4.3 and the supplementary material), with results summarized in Table 1 and qualitative examples in Figure 10. In addition to the full test set, we report performance on subsets ranked by cloud coverage. The benchmark compares against heuristic baselines (LeastCloudy, Mosaicing) and prior zero-shot neural approaches. TEC-TT outperforms all heuristic methods and prior zero-shot neural models on the full test set.
@@
The subset analysis, focusing on the most challenging samples in terms of cloud coverage, further underscores the benefit of temporal priors: heuristic approaches perform adequately in low-cloud cases, but break down under heavy cloud coverage, while TEC-TT maintains high PSNR and SSIM. The hardest 10% of samples correspond to an average cloud coverage of 99%, yet TEC-TT still produces reasonable predictions. Overall, the results demonstrate that the temporal modeling in TEC-TT not only improves compression but also transfers to challenging forecasting tasks.
@@ 6 Conclusion
We introduce and release TerraCodec, a family of learned compression models for Earth observation, pretrained on Sentinel-2 multispectral time series data. Our models outperform classical image and video codecs in rate–distortion, achieving up to an order-of-magnitude reduction at equal quality. Latent Repacking further enables flexible-rate transformer models from a single checkpoint, as demonstrated by FlexTEC. Downstream evaluations show that moderate compression preserves analysis performance, while zero-shot cloud inpainting highlights the strengths of our temporal transformer TEC-TT beyond compression. Overall, TerraCodec establishes a foundation for exploring the benefits of temporal modeling in EO compression and for advancing multispectral learned compression.
@@ 0.B.1 Framework and environment
All models are implemented in PyTorch, using CompressAI [7] for core architectures and entropy bottlenecks, extended to support multispectral inputs, temporal samples, and model-belief analyses. The TEC-TT implementations are based on the Neural compression repository [42]. Experiments are run on NVIDIA A100 GPUs with mixed precision training.
@@ 0.B.2 Image models (TEC-FP, TEC-ELIC)
Checkpoint settings. Hyperparameters $N$ and $M$ denote the channel width of encoder/decoder layers and the latent bottleneck size, respectively. For both codecs, $N$ and $M$ are scaled slightly across the trained $\lambda$ values, as summarized in Table 3. All checkpoints are trained with identical optimization settings (see Sec. 4.1), varying only $N$, $M$, and the rate–distortion trade-off coefficient $\lambda$.
@@
0.B.3 Temporal model (TEC-TT)
We provide additional details on the TEC-TT architecture, including its tokenization strategy and temporal transformer design, following VCT [39].
@@ 0.B.3 Temporal model (TEC-TT)
Training. Training follows the standard uniform-noise quantization surrogate [52, 41], while inference applies hard quantization and arithmetic coding. We train six TEC-TT variants at different rate–distortion trade-offs, controlled by $\lambda\in\{0.4,5.0,20.0,100.0,300.0,700.0\}$, where smaller values enforce higher compression and larger values prioritize reconstruction quality.
@@ Appendix 0.C Latent Repacking and FlexTEC
Implementation differences vs. TEC-TT. FlexTEC is architecturally identical to TECTT except for: (i) the permutation that repacks tokens (Sec. 3.2); (ii) token masking with a learned mask token $m\in\mathbb{R}^{d_{\text{lat}}}$ used during training to replace dropped tokens; and (iii) the masked-rate objective with budget sampling. All other layers, dimensions, and hyperparameters remain unchanged.
@@ 0.E.1 PSNR range sensitivity
In Fig. 11, we plot RD curves for all codecs using PSNR-per-band under different normalization ranges: PSNR 65k (full 16-bit), PSNR 10k (typical Sentinel-2 reflectance 0–10000), and PSNR auto (per-band min–max). We find that PSNR is sensitive to this choice. TerraCodec models (TEC–FP, TEC–ELIC, TEC–TT) remain comparatively stable across ranges, whereas the ranking of classical codecs shifts: JPEG2000 performs best under the PSNR 65k and PSNR 10k ranges, but in the auto setting at higher bitrates it is overtaken by WebP and x265.
@@ 0.E.2 Temporal conditioning effects
Table 6: Effect of inference context $c$ on TEC-TT (trained with 2-frame context, $\lambda{=}5$). We report bits-per-pixel–band–frame (bppbf, $\downarrow$) and PSNR-per-band (65k, $\uparrow$). Context $c\in\{0,1,1{+}1,2\}$ denotes the number of conditioned frames. The P-frames row evaluates only the last two frames (full context) of each 4-frame sequence. Percent changes are relative to c=0c{=}0.
@@ 0.E.3 Flexible-rate behavior
Fig. 14 compares FlexTEC on P-frames (last two frames, full context under c=2c{=}2) versus all frames. The gap is notably larger than for fixed-rate TEC-TT, particularly at low rates. We hypothesize two compounding effects: (i) latent repacking with masked training encourages FlexTEC to concentrate scene-wide, high-utility content into the earliest tokens of bootstrap frames, lowering their cost relative to fixed-rate models; and (ii) for fully conditioned P-frames, the same curriculum distributes information more evenly across tokens, so truncation retains most of the essentials. Together, these effects yield stronger amortization when excluding bootstrap frames, with the benefit most pronounced in the high-compression regime.
Figure 14: FlexTEC RD on P-frames vs. all frames (PSNR-per-band, 65k). The efficiency gains are much more pronounced for the flexible-rate model than for our fixed-rate TEC-TT models. FlexTEC curves use token budgets $K=\{1,2,3,4,5,6,7,8,12,16\}$.

### v3

@@ 1 Introduction
We address these gaps with TerraCodec, a family of neural codecs tailored to multispectral EO and pretrained on Sentinel-2 data. TerraCodec includes efficient image-based models for multispectral inputs: a lightweight Factorized Prior variant (TEC-FP) and an ELIC-based variant (TEC-ELIC) for optimal rate-distortion. We further introduce a Temporal Transformer (TEC-$T$) that captures long-range dependencies across time without relying on hand-crafted motion priors. As shown in Fig. 1, these models offer a significantly better reconstruction than standard codecs at the same compression rate. At equal image quality, they reduce storage size by up to an order of magnitude, TEC-$T$ provides a further reduction for long time series by exploiting temporal structure.
@@ 3.1 TerraCodec
We introduce TerraCodec, a family of learned codecs for EO, including two image codecs: a lightweight factorized prior (TEC-FP) and a stronger space–channel context model (TEC-ELIC), as well as a temporal transformer (TEC-$T$), with a flexible-rate variant (FlexTEC).
@@
Temporal Transformer (TEC-$T$). TEC-$T$ builds on the VCT architecture [39]. We train a transformer to model temporal dependencies of seasonal EO data in latent space, predicting the current frame’s latent distribution from past context. Each frame $\mathbf{x}_{i}$ is encoded to latents $\mathbf{y}_{i}$ and quantized. We partition the current latent into $B$ non-overlapping spatial blocks $\{\hat{\mathbf{y}}_{i,b}\}_{b=1}^{B}$ and the two past latents into overlapping context blocks to increase the transformer’s receptive field. Each block $b$ is flattened into a sequence of $T$ tokens $\{\hat{y}_{i,b,t}\}_{t=1}^{T}$ of channel width $d_{\text{lat}}$. A temporal encoder aggregates the two previous frames into a joint context embedding $z_{\mathrm{joint}}=E(\hat{\mathbf{y}}_{i-2},\hat{\mathbf{y}}_{i-1})$. As shown in Figure 3, within each current block, a masked autoregressive transformer predicts token-wise prior parameters conditioned on already decoded tokens and $z_{\mathrm{joint}}$:
@@
We assume conditional independence across blocks given the context, allowing parallel probability estimation during encoding and parallel block decoding. Causal masking prevents attention to undecoded tokens; see the supplementary material for details. TEC-$T$ uses the same CNN analysis–synthesis transforms as TEC-ELIC. Unlike the original VCT, it is trained end-to-end on the rate–distortion objective without image pretraining, using a $\lambda$-schedule that emphasizes low-rate regimes early. We further adapt TEC-$T$ for flexible-rate scaling, introducing the FlexTEC variant in the next section.
@@ 3.2 Latent Repacking for flexible-rate models
Most neural codecs are trained for a fixed rate–distortion tradeoff. This makes deployment inflexible since achieving different bitrates requires retraining separate models. Our goal is to support variable rates at inference. We introduce Latent Repacking, which redistributes latent channels across tokens, and apply token masking with dynamic rate scaling during training so tokens learn an information-based ordering. Early tokens capture global structure, later ones refine detail. Truncating tokens then lowers bitrate while preserving global content. We demonstrate this by adapting TEC-$T$, where strong priors allow missing tokens to be predicted, making it well-suited for Latent Repacking.
@@
In other words, $t^{\prime}_{u}$ is the $u$-th slice of $k$ consecutive channels of AA, spanning the full spatial field. The procedure is reversible; reapplying slicing and repacking restores the original layout. After repacking, keeping the first $K$ tokens corresponds to $A[:,:,0:K\cdot k],$ the first $K\cdot k$ latent channels at every spatial location.
@@
We apply the approach on TEC-$T$ and introduce the Flexible-Rate TerraCodec (FlexTEC) model. The model applies Latent Repacking and masking inside the temporal transformer after image-wise compression and restores the original layout before image decoding (details in the supplementary material).
@@ 4.1 Pretraining
Temporal codec. TEC-$T$ is trained with a temporal context of two past frames for 300k steps with a global batch size of 24 on four NVIDIA A100 GPUs, requiring about 70 hours. We optimize with AdamW ($\text{lr}\!=\!10^{-4}$, weight decay $10^{-2}$) and employ half-cosine decay schedule to $\eta_{\min}=10^{-6}$ with 15% warmup steps. We sweep six $\lambda$ values for varying rate–distortion settings. For low-rate settings ($\lambda\leq 5.0$), we scale $\lambda$ by 10 during the first 15% of training before annealing back to the target value. Further pretraining details, including FlexTEC, are given in the supplementary material.
@@ 4.3 Downstream tasks
Cloud inpainting. TEC-$T$ captures spatiotemporal priors that can be applied beyond compression. We demonstrate this with zero-shot cloud removal on the AllClear benchmark [66]. Each sample consists of three cloudy observations with masks and one cloud-free target. To apply TEC-$T$, we use the two least cloudy images as a temporal context and extract all cloud-free patches from the least cloudy image as input $x_{0}$. Patches in $x_{0}$ that are covered by clouds are predicted by TEC-$T$. We report PSNR and other metrics, comparing against the official AllClear baselines. Following the benchmark, metrics are computed in auto mode across all bands, not per-band.
@@ 5.1 Rate–distortion
TEC-$T$ improves over the neural image models, although its margin over TEC-ELIC is smaller than the typical video–image gap. While EO sequences differ from natural video, being sampled at daily to seasonal rather than sub-second intervals, the short 4-frame validation setup further departs from typical video settings.
This limits temporal gains because half of the frames are bootstrap frames, which lack full previous context, whereas P-frames are predicted from two past frames. To better understand these effects, we study the role of temporal conditioning and report the RD performance on P-frames only in Figure 6. Without context, TEC-$T$ reduces to an image codec (labeled as image only). For a medium setting ($\lambda=5$), conditioning on one past frame improves compression by 13.6% compared to no context. On P-frames with two previous images, TEC-$T$ achieves a 22.6% rate reduction at equal PSNR, showing that longer EO sequences yield greater efficiency as bootstrap frames are amortized.
@@ 5.2 Flexible rate–distortion
Our TEC-$T$-based FlexTEC model uses Latent Repacking to provide flexible rate-compression from a single checkpoint by transmitting a variable subset of latents and inferring the remainder from the model prior. Figure 8 compares FlexTEC against our fixed-rate models and standard codecs on P-frame compression. While fixed-rate TEC-$T$ models serve as an upper bound—each being optimally fitted for one specific rate–distortion—they require several separately trained models. In contrast, FlexTEC provides several user-controlled rate settings and performs close to or better than TEC-FP, depending on the setting. Further analysis in the supplementary material shows that FlexTEC encodes significant information in bootstrap frames, leading to similar bitrates independent of the token budget. The model then uses this information in the following P-frames to provide efficient rate–distortion settings. Figure 8 provides qualitative examples for different token budgets.
@@ 5.3 Downstream tasks
Model beliefs. Neural codecs rely on learned priors to estimate latent distributions for entropy coding. By decoding these priors into the image space, we obtain predictions that expose the model’s implicit knowledge. Figure 9 shows qualitative examples, with additional results in the supplementary material. We compare priors under three conditions: (a) no information from the current frame (only past context, TEC-$T$), (b) partial information from the current latent (past frames and a subset of tokens for TEC-$T$; subsets of channel groups for TEC-ELIC), and (c) full information of the latent. TEC-FP, lacking a hyperprior, cannot adapt to the latent at hand. Results show that TEC-$T$ already produces plausible forecasts from past frames alone and improves its beliefs when partial information becomes available, yielding the most refined predictions.
Figure 9: Model beliefs obtained by decoding learned priors into the image space. TEC-$T$ is shown with 0 (past context only), 5 and 16 tokens. TEC-ELIC uses limited and full channel-group context.
Cloud inpainting. Building on TEC-$T$’s latent predictions, we evaluate zero-shot cloud removal on the AllClear benchmark [66]. TEC-$T$ is applied without task-specific training (see Sec. 4.3 and the supplementary material), with results summarized in Table 1 and qualitative examples in Figure 10. In addition to the full test set, we report performance on subsets ranked by cloud coverage. The benchmark compares against heuristic baselines (LeastCloudy, Mosaicing) and prior zero-shot neural approaches. TEC-$T$ outperforms all heuristic methods and prior zero-shot neural models on the full test set.
@@
The subset analysis, focusing on the most challenging samples in terms of cloud coverage, further underscores the benefit of temporal priors: heuristic approaches perform adequately in low-cloud cases, but break down under heavy cloud coverage, while TEC-$T$ maintains high PSNR and SSIM. The hardest 10% of samples correspond to an average cloud coverage of 99%, yet TEC-$T$ still produces reasonable predictions. Overall, the results demonstrate that the temporal modeling in TEC-$T$ not only improves compression but also transfers to challenging forecasting tasks.
@@ 6 Conclusion
We introduce and release TerraCodec, a family of learned compression models for Earth observation, pretrained on Sentinel-2 multispectral time series data. Our models outperform classical image and video codecs in rate–distortion, achieving up to an order-of-magnitude reduction at equal quality. Latent Repacking further enables flexible-rate transformer models from a single checkpoint, as demonstrated by FlexTEC. Downstream evaluations show that moderate compression preserves analysis performance, while zero-shot cloud inpainting highlights the strengths of our temporal transformer TEC-$T$ beyond compression. Overall, TerraCodec establishes a foundation for exploring the benefits of temporal modeling in EO compression and for advancing multispectral learned compression.
@@ 0.B.1 Framework and environment
All models are implemented in PyTorch, using CompressAI [7] for core architectures and entropy bottlenecks, extended to support multispectral inputs, temporal samples, and model-belief analyses. The TEC-$T$ implementations are based on the Neural compression repository [42]. Experiments are run on NVIDIA A100 GPUs with mixed precision training.
@@ 0.B.2 Image models (TEC-FP, TEC-ELIC)
Checkpoint settings. Hyperparameters NN and $M$ denote the channel width of encoder/decoder layers and the latent bottleneck size, respectively. For both codecs, NN and $M$ are scaled slightly across the trained $\lambda$ values, as summarized in Table 3. All checkpoints are trained with identical optimization settings (see Sec. 4.1), varying only NN, $M$, and the rate–distortion trade-off coefficient $\lambda$.
@@
0.B.3 Temporal model (TEC-$T$)
We provide additional details on the TEC-$T$ architecture, including its tokenization strategy and temporal transformer design, following VCT [39].
@@ 0.B.3 Temporal model (TEC-$T$)
Training. Training follows the standard uniform-noise quantization surrogate [52, 41], while inference applies hard quantization and arithmetic coding. We train six TEC-$T$ variants at different rate–distortion trade-offs, controlled by $\lambda\in\{0.4,5.0,20.0,100.0,300.0,700.0\}$, where smaller values enforce higher compression and larger values prioritize reconstruction quality.
@@ Appendix 0.C Latent Repacking and FlexTEC
Implementation differences vs. TEC-$T$. FlexTEC is architecturally identical to TECTT except for: (i) the permutation that repacks tokens (Sec. 3.2); (ii) token masking with a learned mask token $m\in\mathbb{R}^{d_{\text{lat}}}$ used during training to replace dropped tokens; and (iii) the masked-rate objective with budget sampling. All other layers, dimensions, and hyperparameters remain unchanged.
@@ 0.E.1 PSNR range sensitivity
In Fig. 11, we plot RD curves for all codecs using PSNR-per-band under different normalization ranges: PSNR 65k (full 16-bit), PSNR 10k (typical Sentinel-2 reflectance 0–10000), and PSNR auto (per-band min–max). We find that PSNR is sensitive to this choice. TerraCodec models (TEC–FP, TEC–ELIC, TEC–$T$) remain comparatively stable across ranges, whereas the ranking of classical codecs shifts: JPEG2000 performs best under the PSNR 65k and PSNR 10k ranges, but in the auto setting at higher bitrates it is overtaken by WebP and x265.
@@ 0.E.2 Temporal conditioning effects
Table 6: Effect of inference context $c$ on TEC-$T$ (trained with 2-frame context, $\lambda{=}5$). We report bits-per-pixel–band–frame (bppbf, $\downarrow$) and PSNR-per-band (65k, $\uparrow$). Context $c\in\{0,1,1{+}1,2\}$ denotes the number of conditioned frames. The P-frames row evaluates only the last two frames (full context) of each 4-frame sequence. Percent changes are relative to c=0c{=}0.
@@ 0.E.3 Flexible-rate behavior
Fig. 14 compares FlexTEC on P-frames (last two frames, full context under c=2c{=}2) versus all frames. The gap is notably larger than for fixed-rate TEC-$T$, particularly at low rates. We hypothesize two compounding effects: (i) latent repacking with masked training encourages FlexTEC to concentrate scene-wide, high-utility content into the earliest tokens of bootstrap frames, lowering their cost relative to fixed-rate models; and (ii) for fully conditioned P-frames, the same curriculum distributes information more evenly across tokens, so truncation retains most of the essentials. Together, these effects yield stronger amortization when excluding bootstrap frames, with the benefit most pronounced in the high-compression regime.
Figure 14: FlexTEC RD on P-frames vs. all frames (PSNR-per-band, 65k). The efficiency gains are much more pronounced for the flexible-rate model than for our fixed-rate TEC-$T$ models. FlexTEC curves use token budgets $K=\{1,2,3,4,5,6,7,8,12,16\}$.